OpenStudy (anonymous):

I am in college algebra and am confused on how to find absolute maximum and minimum on the graphing calculator. the equation is f(x)=x^(2/3)*(x-4)

OpenStudy (anonymous):

\[f \left( x \right)=x ^{\frac{ 2 }{ 3 }}\left( x-4 \right)=x ^{\frac{ 5 }{ 3 }}-4x ^{\frac{ 2 }{ 3 }}\] \[f \prime \left( x \right)=\frac{ 5 }{3 }x ^{\frac{ 2 }{3 }}-4*\frac{ 2 }{3 }x ^{\frac{ -1 }{ 3 }}\] \[f \prime \left( x \right)=\frac{ 5x-8 }{3x ^{\frac{ 1 }{3 }} }\] \[f \prime \left( x \right)=0~ gives \] 5x-8=0, \[x=\frac{ 8 }{5 }\] \[f \prime \prime \left( x \right)=\frac{ 5 }{ 3 }*\frac{ 2 }{ 3 }x ^{\frac{ -1 }{ 3 }}-\frac{ 8 }{ 3 }*\frac{ -1 }{ 3 }x ^{\frac{ -4 }{ 3 }}\] \[f \prime \prime \left( x \right)=\frac{ 10 }{ 9x ^{\frac{ 1 }{ 3 }} }+\frac{ 8 }{ 9x ^{\frac{ 4 }{ 3 } } } \] at x=8/5 find f"(x)

OpenStudy (anonymous):

\[f \prime \prime \left( x \right)=\frac{ 10x+8 }{ 9x ^{\frac{ 4 }{ 3 }} }\] f"(x)>0 at x=8/5 there is minima at x=8/5 ,there is no maxima at x=8/5 put x=8/5 in f(x) and find absolute minima.